OpenStudy (anonymous):

lim h->0(from left) h/sin3h

OpenStudy (dumbcow):

multiply by 3/3 then use identity \[\lim_{x \rightarrow 0}\frac{ax}{\sin ax} = 1\]

OpenStudy (anonymous):

I'm sorry...what do I multiply 3/3 by?

OpenStudy (dumbcow):

the whole expression \[\rightarrow \frac{h}{\sin 3h}*\frac{3}{3} = \frac{1}{3}*\frac{3h}{\sin 3h}\] notice how then you can rearrange the fractions

OpenStudy (anonymous):

oh okay...so I'm moving the 1/3 out front?

OpenStudy (anonymous):

To cancel the 3h/sin3h=1? I think maybe I get it

OpenStudy (dumbcow):

yes because you know that part has a limit of 1 .... then you are just left with the 1/3

OpenStudy (anonymous):

Thank you!!

OpenStudy (dumbcow):

yw